### v3

# Onset of Intermittency and Multiscaling in Active Turbulence

Abstract
Recent results suggest that highly active, chaotic, non-equilibrium states of living fluids might share much in common with high Reynolds number, inertial turbulence. We now show, by using a hydrodynamical model, the onset of intermittency and the consequent multiscaling of Eulerian and Lagrangian structure functions as a function of the bacterial activity. Our results bridge the worlds of low and high Reynolds number flows as well as open up intriguing possibilities of what makes flows intermittent.
Multiscaling
The phenomenon of intermittency is most commonly associated with high-Reynolds-number fluid flows that are turbulent Frisch (1996). An analytical theory of such intermittency has remained elusive because of the formidable challenges posed by the nonlinearity of the Navier-Stokes equations of hydrodynamics  [^1] An analytical theory has been developed for Kraichnan’s model of passive-scalar turbulence Falkovich et al. (2001), in which the passive-scalar equation is linear.. Nevertheless, experiments, observations and numerical simulations of three-dimensional (3D) fluid turbulence (henceforth inertial turbulence) clearly show strong evidence of intermittency Frisch (1996), which manifests itself, inter alia, in the measured deviations of the exponent $\zeta_{p}$ of the order-$p$, equal-time, inertial-range velocity structure functions from the mean-field, dimensional linear $p/3$ result of Kolmogorov Kolmogorov (1991) from 1941 (K41). Specifically, $\zeta_{p}\geq p/3$ for $p<3$ and $\zeta_{p}\leq p/3$ for $p\geq 3$ with $\zeta_{3}=1$ Frisch (1996). Such intermittency or anomalous scaling has also been seen in magnetohydrodynamic Basu et al. (1998), passive-scalar (advected by the Navier-Stokes velocity) Mitra and Pandit (2005) and quantum Boué et al. (2013); Rusaouen et al. (2017); Verma et al. (2023) turbulence. Consequently, in the fluid-turbulence community, a perception has emerged that intermittency and anomalous scaling are a consequence of the high-Reynolds-number fluid turbulence. However, recent results in low-Reynolds systems Liu and I (2012); Qiu et al. (2016); Padhan and Pandit (2023); Singh et al. (2024); Singh and Rosti (2023); Romaguera et al. (2024); Huang (2024), suggest that this is not entirely correct.
A starting point is a reexamination of this perception in dense bacterial suspensions which behave like (living) fluids at approximately zero Reynolds numbers  Dombrowski et al. (2004); Wensink et al. (2012); Dunkel et al. (2013); Marchetti et al. (2013); Ramaswamy (2017); Alert et al. (2022). The emergent fluidised state of such dense suspensions is known to display a variety of dynamical phases James et al. (2021); Mukherjee et al. (2021); Puggioni et al. (2022). Most prominently, for a range of activity $\alpha$, these two-dimensional, low-Reynolds number fluids are marked by chaotic, vortical motion, which is remarkably similar to inertial high-Re turbulence Boffetta and Ecke (2012); Pandit et al. (2017). Hence, the nomenclature active turbulence Wensink et al. (2012) to describe such non-equilibrium states Alert et al. (2022); Bhattacharjee and Kirkpatrick (2022). Not surprisingly, active turbulence has been the subject of several recent studies, some with biological motivations, such as evasion and foraging (Humphries et al., 2010; Volpe and Volpe, 2017), and other investigations of the nature of active turbulence and its relation with fluid turbulence Bratanov et al. (2015); Zanger et al. (2015); James and Wilczek (2018); James et al. (2021); C. P. and Joy (2020); Mukherjee et al. (2021); Kiran et al. (2023); Singh et al. (2022); Gibbon et al. (2023).
There is an additional important reason for us to study this system. The question of multiscaling in active turbulence was examined experimentally a decade ago with contradictory conclusions: The experiments of Liu and I Liu and I (2012) found evidence for the multiscaling of $\zeta_{p}$; by contrast, Wensink et al. Wensink et al. (2012) reported a complete absence of such signatures of intermittency. We resolve this contradiction by showing how active turbulence can display a transition from simple scaling to multiscaling as a function of the activity parameter. This offers a natural explanation for why experiments can lead to varying conclusions regarding intermittency.
Dense, bacterial suspensions follow a hydrodynamic description Wensink et al. (2012); Alert et al. (2020)
$\partial_{t}{\bf u}+\lambda{\bf u}\cdot\nabla{\bf u}$ $=$ $-\nabla p-\Gamma_{0}\nabla^{2}{\bf u}-\Gamma_{2}\nabla^{4}{\bf u}-$ (1)
-- $(\alpha+\beta|{\bf u}|^{2}){\bf u}\,,$
$\nabla\cdot{\bf u}$ $=$ $0\,,$ (2)
where ${\bf u}({\bf x},t)$ is the coarse-grained, incompressible, velocity field of the motile bacterial (active) suspension, with the positive parameter $\lambda$ defining pusher-type bacteria. The chaotic patterns arise from instabilities that arise because $\Gamma_{0},\Gamma_{2}>0$ Swift and Hohenberg (1977); Wensink et al. (2012); Aditi Simha and Ramaswamy (2002); Linkmann et al. (2020); Bratanov et al. (2015). Finally, the Toner-Tu drive Toner and Tu (1995, 1998) is accounted for by the terms with coefficients $\alpha$ and $\beta$. Stability demands $\beta>0$; and the negative activity parameter $\alpha$ injects energy; the more active the suspension the more negative is $\alpha$. The TTSH equation (1) has the same advective nonlinearity as the Navier-Stokes (NS) equation. Hence, it is tempting to ask if there is a range of activity $\alpha$, in which TTSH turbulence has statistical properties that are akin to those of inertial turbulence. In particular, does this activity-driven turbulence possess (approximate) scale-invariance, in a power-law range with a universal spectral scaling exponent, fluctuations, and intermittency that leads to multiscaling?
Figure 1: (a) A plot of the exponent ratios $\zeta_{p}/\zeta_{2}$ (with error bars) obtained from an extended self-similarity (ESS) analysis of the longitudinal structure function versus $p$ for different values of $\alpha\lesssim\alpha_{c}$ (see legend; for $\alpha=-6$ we show exponents from two different simulations with different domain sizes); the dashed black denotes the (linear) scaling $p/2$. In the upper inset we show plots of the second-order (longitudinal) structure functions compensated with the scaling exponent $\zeta_{2}=0.5$ to show our results are consistent with what has been reported earlier for the spectral scaling and the extent our scaling range. In the lower inset, we show plots of $\zeta_{p}/\zeta_{2}$ versus $p$ for $\alpha\gtrsim\alpha_{c}$ and from both sets of simulations (see legend). Clearly, for such mildly active suspensions the scaling exponents are linear with the order $p$. (The error bars are consistent with symbol sizes and not shown explicitly for clarity.). The local slopes $\tfrac{d\log S_{p}}{d\log r}$ vs $r$ for (b) $\alpha=-8$; (c) $\alpha=-6$ and (d) $\alpha=-2.5$. The shaded gray region shows the plateau whose mean value yields the scaling exponents; error-bars follow from the standard deviations.
These questions were partially answered in a recent paper by Mukherjee, et al. Mukherjee et al. (2023) who showed, through a closure analysis complemented by direct numerical simulations, the existence of a critical activity $\alpha_{c}$: For highly active suspensions $\alpha\lesssim\alpha_{c}$, the energy spectrum has a universal activity-independent scaling exponent and the distribution of velocity gradients and increments show distinct non-Gaussian tails. For mildly active suspensions $\alpha\gtrsim\alpha_{c}$, in contrast, the energy spectrum has a non-universal, activity-dependent scaling Bratanov et al. (2015) with Gaussian distributions of velocity gradients and increments Wensink et al. (2012). Furthermore, Kiran, et al. Kiran et al. (2023) have shown that probability distribution functions (PDFs) of temporal increments of the energy, along trajectories of Lagrangian tracers, display wide, non-Gaussian tails. However, the question of the possible multiscaling of both Eulerian and Lagrangian velocity increments, the cornerstone of multiscaling in classical turbulence theory [see, e.g.,  Frisch (1996); Buaria and Sreenivasan (2023a)], has remained unaddressed in active turbulence. We demonstrate that Eulerian and Lagrangian intermittency, in highly active TTSH turbulence, is similar to their inertial-turbulence counterparts.
Our study uses pseudospectral direct numerical simulations of Eqs. (1) on a doubly-periodic domain (James and Wilczek, 2018). We perform different sets of simulations, with domain sizes ranging from $L=2\pi$ to 80, and $1024^{2}$ to $8192^{2}$ collocation points. We use $\Gamma_{0}=0.045$, $\Gamma_{2}=\Gamma_{0}^{3}$, $\beta=0.5$, $\lambda=3.5$ (Wensink et al., 2012; James and Wilczek, 2018; C. P. and Joy, 2020), and the activity parameter $\alpha$ is varied across the transition $\alpha_{c}$.
In analogy with inertial turbulence, we turn our attention to the Eulerian scaling exponents $\zeta_{p}$ of the equal-time, longitudinal, order $p$ structure function $S_{p}(r)\equiv\langle|\Delta{\bf u}\cdot\hat{\bf r}|^{p}\rangle\sim r^{\zeta_{p}}$, where $\Delta{\bf u}={\bf u}({\bf x}+{\bf r})-{\bf u}({\bf x})$. In particular, the second-order exponent $\zeta_{2}$ is trivially associated with the scaling exponent of the energy spectrum $E(k)\sim k^{-\delta}$, for $1\leq\delta\leq 3$, via the Wiener–Khinchin theorem, as $\zeta_{2}=\delta-1$. For $\alpha\lesssim\alpha_{c}$, we know $\delta=3/2$ Mukherjee et al. (2023) and thence $\zeta_{2}=1/2$. We confirm this by showing, in the upper inset Fig. 1 (a), the compensated second order structure function $r^{-\zeta_{2}}S_{2}$ versus $r$ with $\zeta_{2}$ set to 1/2 for $\alpha=-6<\alpha_{c}$. The flat plateau, at values of $r$ consistent with the spectral scaling range Mukherjee et al. (2023), is a confirmation of the self-consistency in our measurements of the structure functions. Furthermore, dimensional analysis suggests that, in the absence of intermittency $\zeta_{p}=p/4$ (the analog of the K41 scaling $p/3$ in inertial turbulence). We now evaluate $S_{p}(r)$ for $p>2$ to uncover the deviations, because of intermittency, from the linear-scaling result $p/4$.
We measure high-order exponents in two ways: First, we use log-log plot of $S_{p}$ vs. $r$ to obtain the exponent from their local slopes, i.e., $\zeta_{p}=\langle\tfrac{d\log S_{p}}{d\log r}\rangle_{r}$. Here, $\langle\cdot\rangle_{r}$ denotes the average over $r$, in the plateau in the upper inset of Fig. 1(a), where the local slope is nearly flat. Second, we also calculate the exponent ratio $\tfrac{\zeta_{p}}{\zeta_{2}}$ by using a local slope analysis in conjunction with the extended-self-similarity (ESS) procedure Benzi et al. (1993); Chakraborty et al. (2010): $\tfrac{\zeta_{p}}{\zeta_{2}}=\langle\tfrac{d\log S_{p}}{d\log S_{2}}\rangle_{r}$. Both these methods yield consistent values of $\zeta_{p}$; understandably, the error bars improve when we employ ESS, which allows us to obtain a scaling range of about a decade as indicated in Figs. 1(b)-(d). The mean exponents are the average values of the plateaus; the error-bars follow from the standard deviation.
In Fig. 1(a) we show plots of $\tfrac{\zeta_{p}}{\zeta_{2}}$ versus $p$, for different values of $\alpha\lesssim\alpha_{c}$, with a thick black line passing through the mean values of these exponents, as a guide to the eye. Within error bars, the exponents for the different values of $\alpha$ overlap; we have checked that this is also true for different resolutions and system sizes. Our measured exponent ratios show significant deviation from the simple linear scaling. This deviation is, reminiscent of the convex, monotonically increasing plots of equal-time exponents in inertial turbulence Frisch (1996); Buaria and Sreenivasan (2023b), is the first, theoretical evidence of multiscaling in active turbulence for values of $\alpha$ beyond $\alpha_{c}$. It substantiates the experimental measurements made by Liu and I Liu and I (2012) as well as the suggestion of an emergent intermittency in measurements of the PDFs of velocity increments and the associated kurtosis Mukherjee et al. (2023).
To study milder activities, we carry out DNSs with $-1\gtrsim\alpha\gtrsim\alpha_{c}$. In this range of $\alpha$, the spectral scaling exponent is given by $\delta=\frac{\tau_{\rm eff}(2\alpha+8\beta E_{\rm tot})}{\lambda}-1\gtrsim 0$, with an effective time-scale $\tau_{\rm eff}$ and total energy $E_{\rm tot}$ of the suspension Rose and Sulem (1978); Boffetta and Ecke (2012); Bratanov et al. (2015); Mukherjee et al. (2023). Hence, there is no simple way to estimate the second-order exponent $\zeta_{2}$ from the spectral exponent $\delta$ for such mildly active suspensions. Nevertheless, it is simple, by using the procedure outlined above, to estimate the $\zeta_{p}$ for such values of $\alpha$. In the lower inset of Fig. 1 (a), we present plots of $\zeta_{p}/\zeta_{2}$ vs. $p$, for $\alpha\gtrsim\alpha_{c}$, which show that $\zeta_{p}/\zeta_{2}$ depends linearly on $p$ as noted for $0\leq p\leq 4$ in Ref. Wensink et al. (2012). Therefore, in this mild activity regime there is no multiscaling.
Figure 2: Probability density functions (PDFs) of the Lagrangian velocity increments $\Delta v_{i}$ (both Cartesian coordinates included) for different time lags $\tilde{\tau}=\tau/(2\sqrt{2}\pi\Gamma_{2}/|\Gamma_{0}|^{2})$ (see legend) and for $\alpha=-6$. Clearly, these distributions are Gaussian (indicated by the dashed line) at large time lags, but show significant non-Gaussian tails as $\tau$ becomes smaller. (The curves are staggered vertically for clarity.)
In inertial turbulence, intermittency manifests itself not only in to the Eulerian measurements, but also, and perhaps more crucially from the point of view of transport, mixing, and dispersion, in Lagrangian studies. Thus, a full appreciation of this activity-induced, emergent intermittency in low-Reynolds number flows requires us to investigate the problem in the Lagrangian framework, where the governing equations for Lagrangian trajectories, $\bf x$, are given by
${}\frac{d{\bf x}}{dt}={\bf v};\leavevmode\nobreak\ {\bf v}({\bf x},t)={\bf u}({\bf x}).$ (3)
We measure, in particular, the Lagrangian exponents $\xi_{p}$ via the Lagrangian structure functions $S_{p}^{\rm L}=\sum_{i}\langle|\Delta v_{i}|^{p}\rangle\sim\tau^{\xi_{p}}$, where $\Delta v_{i}\equiv v_{i}(t+\tau)-v_{i}(t)$, the time lag is $\tau$, and the index $i$ denotes the Cartesian component. We focus first on highly active suspension, so we present results for $\alpha=-6$. [We have checked that our results are robust for all values of $\alpha\lesssim\alpha_{c}$.]
The PDFs of $\Delta v_{i}$ are instructive. For time lags much larger than the instability time scales, i.e., $\tau\gg 2\sqrt{2}\pi\Gamma_{2}/|\Gamma_{0}|^{2}$, these PDFs ought to be Gaussian (irrespective of the value of $\alpha$) as we see indeed in Fig. 2. However, for smaller values of $\tau\lesssim 2\sqrt{2}\pi\Gamma_{2}/|\Gamma_{0}|^{2}$, we show in Fig. 2 that these PDFs develop fat, non-Gaussian tails (for $\alpha\lesssim\alpha_{c}$), a clear indication of Lagrangian intermittency [cf. Refs. Mordant et al. (2001) and Arnèodo et al. (2008) for inertial turbulence]. We characterise these fat tails, below, by direct measurements of $\xi_{p}$.
Before we turn to the actual measurements of the Lagrangian exponents, we recall that there is a simple dimensional argument Biferale et al. (2004); Arnèodo et al. (2008) which leads to a relation between the Lagrangian exponents $\xi_{p}$ and the Eulerian exponents $\zeta_{p}$: If in time $\tau$, the Lagrangian particle moves from ${\bf x}_{1}$ to ${\bf x}_{2}$, with ${\bf r}=|{\bf x}_{2}-{\bf x}_{1}|$ and $r=|{\bf r}|$, and we assume that the corresponding Lagrangian-velocity increment $\Delta v\sim|\Delta{\bf u}\cdot\hat{\bf r}|\sim S_{1}(r)$, then it follows that $\Delta v\sim r^{\zeta_{1}}$. The estimate $r\sim(\Delta v)\tau$ for the distance traveled by the Lagrangian particle , yields $\Delta v\sim\tau^{\frac{\zeta_{1}}{1-\zeta_{1}}}$; this implies that $\xi_{1}=\frac{\zeta_{1}}{1-\zeta_{1}}$. If we assume that simple scaling holds (i.e., there is no intermittency), then $\zeta_{p}=p/4$; and $\xi_{1}=1/3$ and $\xi_{p}=p/3$. (For inertial turbulence, a similar argument gives $\xi_{p}=p/2$ Biferale et al. (2004); Arnèodo et al. (2008).)
To go beyond simple scaling, we must obtain the Lagrangian exponents $\xi_{p}$ from our DNS. Once a statistically steady state is reached, we seed the flow, with $10^{5}$ tracer particles, initially distributed randomly, and obtain their trajectories by integrating Eq. (3). For time marching we use a second-order Runge-Kutta scheme and use bilinear interpolation for $\bf{u}$ at off-grid points.
We first focus on the second-order Lagrangian structure function to confirm $\xi_{2}\simeq 2/3$. Then we use a combination of ESS Benzi et al. (1993); Chakraborty et al. (2010) and local-slope analysis to calculate $\xi_{p}/\xi_{2}$. In the inset of Fig. 3, we show (for $\alpha=-6$) plots of the local slopes $\frac{d\log S^{\rm L}_{p}}{d\log S^{\rm L}_{2}}$ vs. $\tau$. In the limit $\tau\to 0$, the local slopes asymptote $\sim p/2$ beacuse small-scale smoothness yields the Taylor-expansion result $S^{\rm L}_{p}\sim\tau^{p}$. For slightly larger values of $\tau\simeq\mathcal{O}(10^{-1})$, there is a dip in the local slopes, which is familiar from similar measurements in inertial turbulence Arnèodo et al. (2008); Buzzicotti et al. (2016)), where this dip is attributed to the trapping of tracers in vortices. For even larger values of $\tau$, beyond this dip, a new scaling range emerges for $S^{\rm L}_{p}(\tau)$, which corresponds to the scaling range for the Eulerian $S_{p}(r)$ [Fig. 1 (c)], if we relate time and length scales via $\tau\sim r/\Delta v$. In the inset Fig. 3, this range seen appears as a plateau after the dip; the mean value of this plateau and the standard deviation from this mean yield the exponent ratios $\xi_{p}/\xi_{2}$ and their error bars, respectively. If there were no intermittency, we would have found $\xi_{p}=p/3$ and $\xi_{p}/\xi_{2}=p/2$. Our plot the exponent ratios as a function of $p$, in Fig. 3, shows significant deviations from the simple-scaling the $p/2$ line. This clear departure from linear scaling underlines the presence of Lagrangian intermittency in low-Re TTSH turbulence. Furthermore, we find another intriguing similarity with recent results for inertial turbulence Iyer et al. (2020); Buaria and Sreenivasan (2023b): The exponents $\xi_{p}$ saturate to a constant for $p\gtrsim 5$. In inertial turbulence, this saturation of Lagrangian exponents has been attributed to the saturation of the exponents $\zeta^{\perp}_{p}$ of the transverse, equal-time Eulerian structure functions: $S^{\perp}_{p}(r)\equiv\langle|\Delta{\bf u}-(\Delta{\bf u}\cdot\hat{\bf r})\hat{\bf r}|^{p}\rangle\sim r^{\zeta^{\perp}_{p}}$. We uncover a similar correspondence with $\zeta^{\perp}_{p}\approx\xi_{p}$ for $p\gtrsim 5$ (see Fig. 3); these exponent ratios appear to saturate to $\simeq 1.5$. In inertial turbulence, it has been argued such saturation is likely because of slender vortex filaments Frisch (1996); Buaria and Sreenivasan (2023b). We conjecture that the exponent saturation we see, for $\alpha\lesssim\alpha_{c}$, may be related to localised streak-like structures, which have been seen in earlier studies Mukherjee et al. (2021); Singh et al. (2022); Puggioni et al. (2022).
We have also obtained $S^{\rm L}_{p}$ in the mild-activity range $\alpha_{c}<\alpha<0$. We have shown above, the Eulerian structure functions $S_{p}$ display simple scaling in this range. By contrast, the behavior of $S^{\rm L}_{p}$ is more complex inasmuch as they show intermittent deviations from Gaussian statistics, but only at small values of $\tau$. However, the Lagrangian structure functions $S^{\rm L}_{p}$ do not display a well-defined inertial range of scales, of the type we show in Fig. 3 for the high-activity case $\alpha=-6$.
Figure 3: Plots (with error bars) of the Lagrangian exponent ratios $\xi_{p}/\xi_{2}$ and the Eulerian transverse structure function exponent ratios $\zeta^{\perp}_{p}/\zeta^{\perp}_{2}$ (from two different simulations with different domain sizes; see legend) for $\alpha=-6$. The solid black line gives the linear scaling result $p/2$. In the inset we show the local slopes $d\log S^{L}_{p}/d\log S^{L}_{2}$ versus the temporal increment $\tau$ for various orders. The shaded gray region shows the plateau whose mean value yields the scaling exponents; error-bars follow from the standard deviations.
Our Eulerian and Lagrangian analyses of the emergent multiscaling are interesting from the point of view of both turbulence and the hydrodynamics of dense, active suspension. They show that active turbulence in these suspensions shares key fingerprints of inertial turbulence, namely, the multiscaling of both Eulerian and Lagrangian structure function, in the high-activity regime $\alpha\lesssim\alpha_{c}$. This emergent multiscaling, as a function of $\alpha$, also resolves the contradiction in the findings from experiments reported earlier. In the experiments of Wensink et al. Wensink et al. (2012), the variations in the velocity field are small and the typical velocities $\simeq 25\mu m/{\rm sec}$ lead to $\alpha\simeq-1>\alpha_{c}$. Hence, as reported by us in the lower inset of Fig. 1 (a), a simple scaling of $\zeta_{p}$ emerges. The typical large scale velocities in the experiments of Liu and I Liu and I (2012) are similar to those in Ref. Wensink et al. (2012), but show much strong variations in speeds, up to $\simeq 55\mu m/{\rm sec}$, leading to an effective $\alpha\simeq-4\simeq\alpha_{c}$  [^2] We refer the reader to the Supplementary Information in Ref. Wensink et al. (2012) for details on the conversion of experimentally measured velocities to the activity parameter $\alpha$.. Indeed, their multiscaling exponents are consistent with Fig. 1 (a) and, in particular, with our measured $\zeta_{2}$. Significantly, these studies also complement, from a hydrodynamical approach, experimental results Secchi et al. (2016) suggesting a motility-induced transition to intermittent flows in B. Subtilis suspensions.
How is multiscaling directly relevant to the individual active agents which make up our flow? It is tempting to conjecture that the the anomalous diffusion reported earlier for $\alpha\lesssim\alpha_{c}$ Mukherjee et al. (2021); Ariel et al. (2015); Gautam et al. (2024) may well have its underlying roots in the Lagrangian multiscaling we report now. We do know that multiscaling in inertial turbulence has direct bearings on the motion of individual particles and in particular on how aggregates form Bec et al. (2016). Emergent intermittent flow here could lead to similar advantages for individual agents in dense active suspensions.
Beyond possible biological implications of our results, there is a second reason why this work is important. Our work advances significantly the understanding of the dynamics of dense bacterial suspensions by isolating the truly turbulent effects from those stemming from simpler chaotic motion. More intriguingly, and at a broader conceptual framework, this study yet again underlines that intermittency can be an emergent phenomena in flows where the nonlinearity does not, trivially, dominate the viscous damping. Indeed, there is increasing evidence of intermittency emerging in systems which are not turbulent in the classical sense. Examples include flows with modest Reynolds number of $\sim\mathcal{O}(10^{2})$ showing intermittent behaviour characteristic of high Reynolds turbulence Schumacher et al. (2014), self-propelling active droplets with intermittent fluctuations Padhan and Pandit (2023), active matter systems of self-propelled particles, which undergo a glass transition, with an intermittent phase before dynamical arrest Mandal et al. (2020), and perhaps most pertinently, in elastic turbulence Singh et al. (2024); Singh and Rosti (2023). Fundamentally, we present, for the first time, evidence of a critical threshold in the control parameter $\alpha$ which allows a transition from a simple-scaling, non-intermittent flow to a multiscaling, intermittent one. Hence our studies throw up interesting questions and offer a fresh perspective on understanding what causes flows to turn intermittent and help in addressing fundamental questions in high-Reynolds-number inertial turbulence.
Acknowledgements.
R. P. and K. V. K. acknowledge support from the Anusandhan National Research Foundation (ANRF), the Science and Engineering Research Board (SERB), and the National Supercomputing Mission (NSM), India, for support, and the Supercomputer Education and Research Centre (IISc), for computational resources. KK and SSR are indebted to S. Mukherjee for his tutelage on understanding active turbulence as well as developing the DNS code at ICTS. The CPU/MPI simulations were performed on the ICTS clusters Tetris and Contra. The GPU simulations were performed on the CUDA A100 at Departement of Physics, IISc. KVK and SSR thank Rahul Singh for careful reading of the manuscript. SSR acknowledges SERB-DST (India) projects STR/2021/000023 and CRG/2021/002766 for financial support. KK and SSR acknowledges the support of the DAE, Govt. of India, under project no. 12-R&D-TFR-5.10-1100 and project no. RTI4001. AG acknowledges SERB-DST (India) projects MTR/2022/000232, CRG/2023/007056-G for financial support. This research was supported in part by the International Centre for Theoretical Sciences (ICTS) for the discussion meeting- Turbulence and Vortex dynamics in 2D quantum fluids, code: ICTS/QUFLU2024/02.